OpenStudy (anonymous):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) x + y = 17 y = x + 7

OpenStudy (anonymous):

The question is trying to trick you. Don't solve it!!!

OpenStudy (nikato):

So u can substitute the second equation into the first x+x+7=17 And solve for x

OpenStudy (anonymous):

10

OpenStudy (anonymous):

?

OpenStudy (nikato):

No, subtract 7 on both sides. And combine like terms. 2x=10 so x=?

OpenStudy (anonymous):

5

OpenStudy (nikato):

Correct. So x=5 Now plug that value to any one of the equation, so y=5+7 so y=?

OpenStudy (anonymous):

12?

OpenStudy (nikato):

So so x=5 and y=12 So ur answer is (5,12)

OpenStudy (anonymous):

ok thankyou!

OpenStudy (nikato):

Your welcome

OpenStudy (anonymous):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) y = 3x + 5 x = -3y - 5

OpenStudy (nikato):

U need help with this too?

OpenStudy (anonymous):

yes